OpenStudy (anonymous):

Medals given! Pleease help! Solve for x. -ax + 2b > 8 x < the quantity negative 2 times b plus 8 all over negative a x > the quantity negative 2 times b plus 8 all over negative a x < the quantity 2 times b minus 8 all over negative a x > the quantity negative 2 times b minus 8 all over a

OpenStudy (amoodarya):

a>0 ? or a<0

OpenStudy (anonymous):

These are the answer choices. A.x < the quantity negative 2 times b plus 8 all over negative a B. x > the quantity negative 2 times b plus 8 all over negative a C. x < the quantity 2 times b minus 8 all over negative a D. x > the quantity negative 2 times b minus 8 all over a

OpenStudy (amoodarya):

-ax + 2b > 8 suppose a<0 then -ax>8-2b -a>0 so divide by -a x>(8-2b)/(-a) for a<0

OpenStudy (anonymous):

would the answer be B?

OpenStudy (amoodarya):

yeah

OpenStudy (sleepyjess):

−ax+2b>8 −ax>−2b+8 −ax>8−2b −ax−1a<8−1a−2b−1a ax1a<8−1a−2b−1a axa<8−1a−2b−1a x<8−1a−2b−1a x<−81a−2b−1a x<−8a−2b−1a x<−8a+2b1a x<−8a+2ba x<2ba−8a

OpenStudy (anonymous):

−ax+2b>8 −ax>−2b+8 −ax>8−2b −ax−1a<8−1a−2b−1a ax1a<8−1a−2b−1a axa<8−1a−2b−1a x<8−1a−2b−1a x<−81a−2b−1a x<−8a−2b−1a x<−8a+2b1a x<−8a+2ba x<2ba−8a i got the same answer as some other guy but.... whatever

OpenStudy (sleepyjess):

Should be 2b-8 over a, but I don't know how to do fractions on here, so I say the answer is C

OpenStudy (anonymous):

Is the answer C or B?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I say C

OpenStudy (anonymous):

i got -ax+2b>8 x>8-2b/-a

OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

Well...

OpenStudy (anonymous):

-Thinks give me a sec-

OpenStudy (anonymous):

thanks <3

OpenStudy (anonymous):

z - 2 = -4... i think

OpenStudy (anonymous):

if i helped give me a medal....

OpenStudy (anonymous):

These are the answer choices. x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

its x -2 = -4

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x + 5 > 9 and 5x less than or greater to 30. A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right. A number line with an open circle on 2, a closed circle on 6, and shading in between. A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on 2, an open circle on 6, and shading in between.

OpenStudy (anonymous):

-Reads- -Hmmmmm-

OpenStudy (sleepyjess):

See attachment

OpenStudy (anonymous):

Srry im kinda stomped on this 1....

OpenStudy (anonymous):

2x + 5 > 9 2x + 5-5 > 9-5 2x > 4 2x/2 > 4/2 x > 2

OpenStudy (anonymous):

So 2x + 5 > 9 turns into x > 2

OpenStudy (anonymous):

5x <= 30 5x/5 <= 30/5 x <= 6 ------------- So 5x <= 30 turns into x <= 6

OpenStudy (anonymous):

this means 2x + 5 > 9 and 5x less than or greater to 30 is the same as x > 2 and x <= 6

OpenStudy (anonymous):

assuming that the "less than or greater to" is actually "less than or equal to"

OpenStudy (anonymous):

theres your answer...?

OpenStudy (anonymous):

@DatSmartGuy yeah

OpenStudy (anonymous):

A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right. A number line with an open circle on 2, a closed circle on 6, and shading in between. A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on 2, an open circle on 6, and shading in between. which is the best answer? :3

OpenStudy (anonymous):

Where's your answer? @DatSmartGuy

OpenStudy (anonymous):

@Trojan i didn't have one....

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

@MyChem Did I help you?

OpenStudy (anonymous):

you all helped but I'm not sure what the correct answer was for the last question I posted.

OpenStudy (anonymous):

Which was...?

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x + 5 > 9 and 5x less than or greater to 30. A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right. A number line with an open circle on 2, a closed circle on 6, and shading in between. A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on 2, an open circle on 6, and shading in between.

OpenStudy (anonymous):

Oh...................................

OpenStudy (anonymous):

yah i just need the answer so I can turn it in

OpenStudy (anonymous):

Well chem the answer is.... Let me just do i'll give it to you....

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

Its a number line with a closed circle on 2, an open circle on 6, and shading in between.

OpenStudy (anonymous):

Does this mean I get a medal or fan...?

OpenStudy (anonymous):

a medal! ;D

OpenStudy (anonymous):

could you help with one more pleasee?

OpenStudy (anonymous):

The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units? No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64. No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32. Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64. Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32.